It says 85 is wrong and I am confused
Check my work mode : This shows what is correct or incorrect for the work you have completed so for. t does not indices Problem 02.033 - Calculation of unknown gradient For the cash flow revenues shown below. find the value of G that makes the equivalent annual worth in years 1 through 7 equal to $500. The interest rate is 7% per year. sh flow, $ ash flow, $ 25 25+G 25 +2G 225 +3G 25 +4G 225+50G The value of G is $ 85 0 AHW 1.pd AHW 1 (1) pdf
Answer #1
Year Cash flow PVIF@7% Present value of cash flow
1 225 0.9346 210.285
2 225 + G 0.8734 196.515 + 0.8734*G
3 225 + 2G 0.8163 183.6675 + 1.6326*G
4 225 +3G 0.7629 171.6525 + 2.2887*G
5 225 + 4G 0.7130 160.425 + 2.852*G
6 225 + 5G 0.6663 149.9175 + 3.3315*G
7 225 + 6G 0.6227 140.1075 + 3.7362*G
Total 1213.135 + 14.7144*G

Equivalent annual worth cash flow = Present value of cash flow / (P/A, 7%, 7)

=1213.135 + 14.7144*G / 5.389

= 225.11319 + 2.3045*G

Equivalent annual worth cash flow = $500

500 = 225.11319 + 2.3045*G

274.8868 = 2.3045*G

G = 119.28